proof identity (1-tanA)/(1+tanA)=(cos2A)/(1+sin2A)
1
Expert's answer
2012-11-05T10:35:15-0500
(1-tanA)/(1+tanA)=(1-sinA/cosA)/(1+sinA/cosA)=(cosA-sinA)/(cosA+sinA)=(cos^2A-sin^2A)/(cosA+sina)^2=cos2A/(sin^2A+cos^2A+2sinAcosA